OpenStudy (anonymous):

Factor by Difference of Two Squares. 1/16 - y^2

OpenStudy (saifoo.khan):

\[\sqrt{\frac{1}{16}} - \sqrt{y^2} = (\frac14 - y)(\frac14 + y)\]

OpenStudy (anonymous):

okay thank you!